OpenStudy (anonymous):

Which pair of functions are inverses?a. f(x) = 2, g(x) = -2 c. f(x) = -x + 3, g(x) = x - 3 b. f(x) = x2 , g(x) = x - 3 d. f(x) = x + 1, g(x) = 2x - 2

OpenStudy (anonymous):

a. f(x) = 2, g(x) = -2 c. f(x) = -x + 3, g(x) = x - 3 b. f(x) = x2 , g(x) = x - 3 d. f(x) = x + 1, g(x) = 2x - 2

hartnn (hartnn):

you know how to find inverse... which one troubles you ?

OpenStudy (anonymous):

I dont see where I plug it in

OpenStudy (anonymous):

OH I FEEL STUPID

OpenStudy (anonymous):

neverrrrrmindddddd

OpenStudy (anonymous):

you would have swap the variables and solve for y

hartnn (hartnn):

so, you did figure out all of them ?

OpenStudy (anonymous):

dont feel stupid :P

OpenStudy (anonymous):

a quick way would be to plug in a test number but then it could also be the danger that the test point will work but it wont be an inverse function